Messages 0xx


01: Epi is working (Robin King)
02: Epi reservation (Nick Lindert)
03: Re: Epi reservation (Robin King)
04: reservation cancellation (Robin King)
05: Epi Monday (Nick Lindert)
06: reservatin' for epi (Aaron Partridge)
07: Low temp epi growth (Nick Lindert)
08: Epi shutdown (Robin King)
09: Monday Epi Reservation (Nick Lindert)
10: Epi use today (Nick Lindert)
11: epi Tues night (Nick Lindert)
12: Re: epi Tues night (Keith Gaul)
13: epi tonite (Nick Lindert)
14: Tube condition (Keith Gaul)
15: epi tonite (Nick Lindert)
16: OK to run (Robin King)
17: thurs nite epi reservation (Nick Lindert)
18: Re: thurs nite epi reservation (Robin King)
19: Re: thurs nite epi reservation (Nick Lindert)
20: epi use on Monday (Nick Lindert)
21: epi wednesday (Nick Lindert)
22: HCl Problem - (Keith Gaul)
23: Epi ready for use (Keith Gaul)
24: Epi looks good (Nick Lindert)
25: epi weirdness (Nick Lindert)
26: germane source (Dan Aubertine)
27: please respond re: doped epi (Robin King)
28: tube change (Ray Seymour)
29: Tube change on 3/7/01 (Ray Seymour)
30: system update (Ray Seymour)
31: wafer wand (Ray Seymour)
32: shutdown progress (Ray Seymour)
33: EPI repair status (Dick Crane)
34: Toxic gas shutdown Aug. 6 (Dick Crane)
35: Toxic gas tests complete (Dick Crane)
36: System status (Dick Crane)
37: HCL Leak Check (Ray Seymour)
38: diclor line (Ray Seymour)
39: HCL cylinder (Ray Seymour)
40: HCL change (Ray Seymour)
41: Repair update (Dick Crane)
42: epi repair date (Dick Crane)
43: epi backup service (Dick Crane)
44: epi repair update (Dick Crane)
45: EPI testing (Dick Crane)
46: Re: EPI testing (maurice stevens)
47: Epi Free Wed Afternoon (Michael S. Bartsch)
48: cancelled afternoon res. (Rob Norris Candler)
49: cancelled res. (Rob Norris Candler)
50: Cancel res for Monday 3/4 (hlkung.stanford.edu)
51: EPI temp uniformity repair (Dick Crane)
52: Canceling EPI time Friday 10-4 (Helen L. Kung)
53: EPI update (Dick Crane)
54: HCl ready to use (Dick Crane)
55: Wafer stuck (?) in epi (Michael S. Bartsch)
56: Nevermind... (Michael S. Bartsch)
57: Shut Down (Michael S. Bartsch)
58: Epi Up (Michael S. Bartsch)
59: wafer left in left load lock (hlkung.stanford.edu)
60: SiGe graded buffer layers (Dan Aubertine)
61: Re: DCS] (Mike Dickey)
62: Re: Shutdown epi 2002-03-28 12:13:02: DCS sensor in gas vault alarming] (Ray Seymour)
63: Re: Shutdown epi 2002-03-28 12:13:02: DCS sensor in gas vault alarming] (Maurice Stevens)
64: epi down/ leak checking (Maurice Stevens)
65: done w/epi for today (Rob Norris Candler)
66: Sorry for the very late notice, but I won't be needing my epi time (Rob Norris Candler)
67: Finished epi early, afternoon is now available (Rob Norris Candler)
68: not using my time (auber.stanford.edu)
69: epi free today -- wafers not ready. sorry for late notice (Aaron Partridge)
70: Loss of house vacuum on Monday6-17-02 (Ted Berg)
71: Re: Loss of house vacuum on Monday6-17-02] (Ted Berg)
72: Ge & SiGe (Gary Yama)
73: Low Temp SiGe growth? (Dan Aubertine)
74: HCL Change (Ray Seymour)
75: HCL (Ray Seymour)
76: EPI Progress report (Dick Crane)
77: Germane Source (auber.stanford.edu)
78: Reservation cancellation (Ted Berg)
79: Tube change (Ted Berg)
80: Tube replacement (Ted Berg)
81: Epi up (Maurice Stevens)
82: Re: outside source for SiGe (30% Ge) on Si? (Jim McVittie)
83: Epi Process Poll (Jim McVittie)
84: Sump Pump is dead for epi (Ted Berg)
85: Re: Comment epi 2003-01-09 13:15:35: HCl bottle changed friday/down over weekend] (Ray Seymour)
86: HCL change (Mike Dickey)
88: Re: cancelled res. Monday 1300-1900 (Benjamin Chui)
89: This weekend...No 1% PH3 (Maurice Stevens)
90: epi available monday 4/14 (Rob Norris Candler)
91: dropped 4/29 reservation (Rob Norris Candler)
92: epi shut down for tube clean (Dick Crane)
93: Epi Tube Pull on Tues (Maurice Stevens)
94: lightly doped N type Si recipe (Yahong Yao)
95: Status if Epi (Ted Berg)
96: epi staus...yellow (Maurice Stevens)
97: EPI STATUS...YELLOW (Maurice Stevens)
98: grabbing of epi-time (Steffen Walker)
99: epi open 18:00-- today (Yahong Yao)